OpenStudy (anonymous):

Solve the equation you found for “the sum of x and 13 is 26”. Show your work please!

OpenStudy (anonymous):

x+13=26 subtract 13 from both sides x=13

OpenStudy (anonymous):

thanx:)

OpenStudy (anonymous):

welcome :D

OpenStudy (anonymous):

you wanna help with more cause i got a ton! it alright if u say no btw

OpenStudy (anonymous):

sure lovely :D

OpenStudy (anonymous):

kay! What is the mathematical expression for “the $100 earned was split between all the people and each got $12.50”? Answer 100 – p = 12.50 100/p = 12.50 100p = 12.50 100p + 12.50 a?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

oh... thanx:)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

ya girl

OpenStudy (anonymous):

A certain amount of money was split between 8 people, and each person got $14.50. How much money was divided?

OpenStudy (anonymous):

?/p=14.50 ?/8=14.50 multiply 8 both sides ?= (answer here)

OpenStudy (anonymous):

i aint got a calculator:9 what should i do???

OpenStudy (anonymous):

116

OpenStudy (anonymous):

ok thanx element.i feel like im using u for answers:(

OpenStudy (anonymous):

Its alright i like helping... sometimes

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

:D ask anytime im on and ill help

OpenStudy (anonymous):

kay ill stop getting answers now

OpenStudy (anonymous):

ok i mean ill still help

OpenStudy (anonymous):

ur choice element

OpenStudy (anonymous):

:D im with you all da way

OpenStudy (anonymous):

okay then:) What is the mathematical expression for “double my age then increase it by 12, the result is 62”. Answer 2a + 12 = 62 2(a + 12) = 62 2 + a + 12 = 62 2a – 12 = 62

OpenStudy (anonymous):

A

OpenStudy (anonymous):

:) i have ten. this is five. still go?

OpenStudy (anonymous):

ya girl im doin this for you :D

OpenStudy (anonymous):

kay! then im go:) Using the equation you found for “double my age then increase it by 12, the result is 62” and the ages 23, 25 and 27, showing work for all three values, tell which is a solution and which is not a solution to the equation.

OpenStudy (anonymous):

so im solving for the persons age?

OpenStudy (anonymous):

yes but if u want i can try to help

OpenStudy (anonymous):

nah it good... one sec

OpenStudy (anonymous):

1. 2a+12=62 2(23)+12=62 46+12=62 58=62 no 23 is not the answer 2. 2a+12=62 2(25)+12=62 50+12=62 62=62 YES 25 is the right answer 3. 2a+12=62 2(27)+12=62 54+12=62 66=62 no 27 is incorrect

OpenStudy (anonymous):

What is the mathematical expression for “7 less than the product of 3 and x is equal to x increased by 15”? Answer 7 – 3x = x + 15 7 – 3x = x – 15 3x – 7 = x – 15 3x – 7 = x + 15

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ur so sweet!! thanx

OpenStudy (anonymous):

thx girl and your welcome :D

OpenStudy (anonymous):

okay ill do the rest now:)

OpenStudy (anonymous):

if you still need help im here

OpenStudy (anonymous):

i feel bad cause ur doin everything!:(

OpenStudy (anonymous):

nah its ok

OpenStudy (anonymous):

u sure??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Using the equation you found for “7 less than the product of 3 and x is equal to x increased by 15” and the values x = 8, 11, 15, showing work for all three values, tell which is a solution and which is not a solution to the equation.

OpenStudy (anonymous):

1. 3x – 7 = x + 15 3(8)-7=8+15 24-7=23 17=23 no 8 is not correct 2. 3x – 7 = x + 15 3(11)-7=11+15 33-7=26 26=26 YES 11 is correct 3. 3x – 7 = x + 15 3(15)-7=15+15 45-7=30 38=30 no 15 is not correct

OpenStudy (anonymous):

thanx:)

OpenStudy (anonymous):

yw anymore?

OpenStudy (anonymous):

no im good:)

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

ya i got this! ur probley tired anyway:)

OpenStudy (anonymous):

nah its fine

OpenStudy (anonymous):

JON! DONT BE SO HARDHEADED! ILL BE FINE:)

OpenStudy (anonymous):

NOPE IM HELPING YOU

OpenStudy (anonymous):

-sigh- next question...

OpenStudy (anonymous):

YAY MATH!!!

OpenStudy (anonymous):

Write your own English expression and the corresponding mathematical equation. Once you have the mathematical equation, test three values in the equation. Show all steps. Was any one of the values you chose the solution? hard one:(thanx jon:)

OpenStudy (anonymous):

ok josie this is a trial and error type question... i think it is asking you if you were correct in guessing one of the 3 numbers from the questions above... ex. the last one choices: 8,11,15... so if you had chosen one at first and gotten it right, you would explain why you guessed it, show the work, and write if your guess was correct or not

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

sorry if it got confusing

OpenStudy (anonymous):

thanx:) its fine

OpenStudy (anonymous):

yw anything else?

OpenStudy (anonymous):

no im done for the night with math! im closing the question

OpenStudy (anonymous):

there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im just gonna message and chat now jon

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

you should at least give me a medal lol... im greedy

OpenStudy (anonymous):

ONLY CAUSE I LOVE YOU!

OpenStudy (anonymous):

I LOVE YOU TOO

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

:)